# Gate-tunable nonlocal Josephson effect through magnetic van der Waals bilayers

### v1

@@ Abstract
It is well-known that the proximity effect at superconductor/ferromagnet (S/F) interfaces produces damped oscillatory behavior of the Cooper pair wave function within the ferromagnetic regions, which is analogous to the inhomogeneous Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) superconductivity. It is often called mesoscopic FFLO state and gives rise to $0-\pi$-transitions in S/F/S Josephson junctions. This paper offers an analysis of the proximity effect at interfaces between superconductors and magnetic van der Waals (vdW) bilayers. The specific feature of the proximity effect in the vdW bilayer systems is the presence of non-local Cooper pairs. We predict that the mesoscopic FFLO state formed by such pairs is sensitive to the difference between on-site energies of the monolayers composing the bilayer and, thus, can be controlled by applying a gating potential to one of the monolayers. This opens the possibility of implementing gate-controlled $0-\pi$ transitions in Josephson junctions through the magnetic vdW bilayer weak links.
@@ I Introduction
The proximity effect between different materials modifies their properties near the interface. In the case of a superconductor/normal metal (S/N) interface, the Cooper pairs can penetrate into the normal metal at a typical distance of the order of the normal coherence length $\xi_{N}$. In the diffusive case this length is of the order of $\xi_{N}\sim\sqrt{D/2\pi T}$, where $D$ is the diffusion constant and $T$ is temperature. In the ballistic case $\xi_{N}\sim v_{F}/2\pi T$, where $v_{F}$ is the electron Fermi velocity in the normal metal. At superconductor/ferromagnet (S/F) interfaces the singlet Cooper pairs penetrating into the ferromagnetic region are partially converted into triplet ones and acquire finite momentum. This is caused by the presence of the exchange field $h$ in the ferromagnet. Because of this field, the energy of a spin-up electron decreases by $h$, while the energy of a spin-down electron increases by $h$. To compensate this energy difference, the kinetic energy of the spin-up (down) electron increases (decreases), what gives rise to the difference of the absolute values of their momenta [1, 2, 3]. Thus the pair acquires a finite momentum. Due to the finite momentum the decay of the singlet and triplet pairs into the ferromagnet is superimposed by oscillations. In ballistic case the spatial scale of this oscillations is $\sim v_{F}/2h$ if $h\gg T$ [4, 5]. In diffusive case the penetration length and the period of oscillations are of the same order of the magnetic coherence length $\xi_{F}\sim\sqrt{D/h}\ll\xi_{N}$. Presence of inhomogeneous magnetization in the interlayer region essentially modifies the triplet proximity effect in ferromagnetic interlayers of JJs giving rise to the long-range range triplet component, which does not oscillate and decays at $\xi_{N}$ [3, 6, 7].
@@
The oscillations of the proximity-induced singlet superconducting correlations in the ferromagnet lead to $0-\pi$ transitions of the ground state superconducting phase difference in superconductor/ferromagnet/superconductor Josephson junctions (S/F/S JJs) upon varying the length of the ferromagnetic weak link [8, 2]. Spatial oscillations of proximity-induced triplet superconductivity were also predicted at superconductor/antiferromagnet (S/AF) interfaces [9]. Yet the physical mechanism is different from S/F heterostructures and does not lead to oscillations of the singlet component. Thus, it does not result in $0-\pi$ transitions in S/AF/S JJs. At the same time, $0-\pi$ transitions in S/AF/S JJs upon atomic variations of the AF interlayer length were predicted [10, 11, 12]. The physical reason of such $0-\pi$ transitions is related to the reconstruction of the Andreev bound states spectra inside the interlayer and thus atomically sharp S/AF interfaces are required to observe the effect.
@@
JJs with the $\pi$ ground state phase difference ($\pi$-JJs) are implemented as $\pi$-phase shifters in superconducting and quantum electronics [13, 14, 15]. The broad application prospects motivated active search for novel principles and possibilities of external control of the switching between the 00 and $\pi$ ground states. Various suggestions of externally controlled $0-\pi$ transitions in Josephson junctions include the temperature induced $0-\pi$ transitions [16], $0-\pi$ transitions induced by the electrostatic gating [17, 18] in quantum dot JJs, spin-independent [19, 20, 21, 22, 23, 24, 25, 26] and spin-dependent [27, 28] nonequilibrium quasiparticle distribution, via the manipulation by the mutual orientation of the ferromagnets of a spin-valve inside a JJ [29] and via the manipulation by the exchange field orientation in S/F/S JJs with spin-orbit coupling [30, 31]. In a two-dimensional electron gas (2DEG) $0-\pi$ transitions can also be controlled by an applied magnetic field or gating [32]. The supercurrent-induced long-range triplet component in Josephson junctions [33] provides a possibility of a supercurrent-controllable $0-\pi$ transition [34].
@@
In this paper we consider penetration of proximity-induced superconducting correlations into bilayer magnetic vdW materials or heterostructures. It is demonstrated that at low electron transparency (interlayer hopping) between the interacting vdW layers composing the bilayer the proximity-induced Cooper pairs can be divided into local and nonlocal ones. The penetration of the nonlocal pairs into nonsuperconducting materials substantially differs from the penetration of the local pairs. In particular, the length of spatial oscillations of such pairs can be tuned by applying the gating potential to one of the vdW layers. It provides a way to identify the contribution of such pairs into the Josephson current and to generate $0-\pi$ transitions via gating. It is worth noting that the possibility to control $0-\pi$ transitions via gating results from the unique sensitivity of spatial behavior of the nonlocal pairs to this parameter and is not related to any tuning of the electron density as in experiments with quantum dots and 2DEG [17, 18, 32]. Further we demonstrate that at higher values of the interlayer hopping elements there is a region of strong hybridization between the local and nonlocal pairs, where the physical properties of all pairs differ from the conventional bulk behavior and the corresponding oscillation periods can be controlled by gating, what again gives rise to unusual and tunable behavior of the Josephson current. We investigate and compare manifestations of the nonlocal proximity-induced superconducting pairs in magnetic bilayers with antiferromagnetic, ferrimagnetic and ferromagnetic interlayer ordering.
The paper is organized as follows. In Sec. II the system under consideration is described. Sec. III is devoted to investigation and physical description of different proximity-induced Cooper pairs and their penetration into non-superconducting bilayers. In Sec. IV results for the Josephson current through magnetic vdW bilayers are presented: in Sec. IV.1 the contribution of the nonlocal pairs in JJs with AF bilayer weak links is discussed; Sec. IV.2 demonstrates how these nonlocal pair contribution gives rises to $0-\pi$ transitions induced by gating and in Sec. IV.3 we consider how the nonlocal pairs manifest themselves in JJs with ferromagnet/normal metal interlayers. Our conclusions are formulated in Sec. V. Appendix contains technical details of derivation of the quasiclassical Eilenberger equation for description of bilayer vdW heterostructures.
@@ II System and model
Figure 1: Sketch of the considered JJ. Superconducting sites are blue, red and orange layers represent upper ferromagnetic layer (F1) and bottom ferromagnetic layer (F2), respectively. The interlayer magnetic ordering can be of different types: AF, F or ferrimagnetic, see text for more details. Different hopping elements are also shown.
System under consideration is shown in Fig. 1. It represents a planar JJ made of layered superconducting leads and a bilayer magnetic weak link. The system is modelled by the following tight-binding Hamiltonian on a square lattice:
@@
Here $c_{\bm{i},\sigma}$ is an annihilation operator for electrons at site $\bm{i}$ and for spin $\sigma=\uparrow,\downarrow$. $\mu_{\bm{i}}$ is the on-site energy, which is assumed to be $\mu_{\bm{i}}=\mu_{1,2}$ at sites belonging to the upper (F1) and lower (F2) ferromagnetic layers, respectively, and $\mu_{\bm{i}}=\mu_{S}$ at the sites belonging to the superconducting leads. The on-site energy determines the filling factors of the conduction band of the materials and in the case of isolated layers correspond to their chemical potentials. We assume only nearest-neighbor hopping with $t_{\bm{i}\bm{j}}=t_{x}^{F}$ being the intralayer hopping element for both ferromagnetic layers and $t_{\bm{i}\bm{j}}=t_{z}^{F}$ being the hopping element between F1 and F2 layers. Anisotropic intralayer (interlayer) hopping elements between nearest superconducting sites are denoted by $t_{\bm{i}\bm{j}}=t_{x(z)}^{S}$ and $t_{\bm{i}\bm{j}}=t^{SF}$ is the hopping element between the upper ferromagnetic layer and the superconducting leads. $\langle\bm{i}\bm{j}\rangle$ means summation over nearest neighbors. $h.c.$ means hermitian conjugate.
$\bm{h}_{\bm{i}}$ is the exchange field in the ferromagnetic layers, which is equal to $\bm{h}_{1,2}$ in the F1 and F2 layers, respectively and is assumed to be spatially homogeneous in the plane of each layer. In superconducting leads $\bm{h}_{\bm{i}}=0$. Below we consider different physical cases: $\bm{h}_{1}=\bm{h}_{2}$ (ferromagnetic interlayer ordering), $\bm{h}_{1}=-\bm{h}_{2}$ (antiferromagnetic interlayer ordering), $\bm{h}_{1}$ is opposite to $\bm{h}_{2}$ and $|\bm{h}_{1}|\neq|\bm{h}_{2}|$ (ferrimagnetic interlayer ordering). The case $\bm{h}_{1}\neq 0$ and $\bm{h}_{2}=0$ modelling a bilayer consisting of a ferromagnetic and a paramagnetic layers is also considered. Such a bilayer structure can be implemented by exfoliation of vdW magnetic materials [46, 47, 48] or using chemical vapor deposition (CVD) processes. Thin films of some intermetallic materials, such as $\mathrm{GdIr_{2}Si_{2}}$ or $\mathrm{TbRh_{2}Si_{2}}$, seem to be good candidates [49, 50]. Also the bilayers can be assembled from magnetic and nonmagnetic monolayers of different vdW materials [40], possibly with an insulating vdW layer between them in order to exclude the direct exchange interaction and to control the magnetization orientations in the layers by the applied magnetic field.
@@ III Penetration of nonlocal correlations into nonsuperconducting regions
Here we consider a non-superconducting bilayer, which has an interface with a superconductor and investigate the proximity penetration of the Cooper pairs from the superconductor into the depth of the bilayer. Details of the geometry of the interface do not matter in this consideration. The superconductor can be located on top of a part of the bilayer (a half of the sketch presented in Fig. 1 without the second lead) or the superconductor can have a contact with both non-superconducting layers. These details only influence the amplitudes of different types of the Cooper pairs induced in the bilayer by proximity effect, yet do not affect their spatial behavior inside the bilayer, penetration depth and oscillating period. The problem can be studied analytically in the framework of the Eilenberger equation for the quasiclassical Green’s function describing electronic behavior in the bilayer. The quasiclassical Green’s function $\check{g}(\bm{R},\bm{p}_{F})$ in the two-layer formalism is a $8\times 8$ matrix in the direct product of spin, particle-hole and layer spaces. It depends on the center of mass coordinate $\bm{R}=(\bm{i}+\bm{j})/2$ of the paired electrons with momenta $\bm{p}_{F}$ and $-\bm{p}_{F}$ at the Fermi surface and the direction of the trajectory determined by $\bm{p}_{F}$. In the non-superconducting bilayer with zero order parameter $\Delta_{\bm{i}}=0$ in both layers it obeys the following generalized Eilenberger equation (the details of the derivation of this equation are provided in Appendix):

### v2

@@ Abstract
It is well-known that the proximity effect at superconductor/ferromagnet (S/F) interfaces produces damped oscillatory behavior of the Cooper pair wave function within the ferromagnetic regions, which is analogous to the inhomogeneous Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) superconductivity. It is often called the mesoscopic FFLO state and gives rise to $0-\pi$-transitions in S/F/S Josephson junctions. This paper offers an analysis of the proximity effect at interfaces between superconductors and magnetic van der Waals (vdW) bilayers. The specific feature of the proximity effect in the vdW bilayer systems is the presence of non-local Cooper pairs. We predict that the mesoscopic FFLO state formed by such pairs is sensitive to the difference between on-site energies of the monolayers composing the bilayer and, thus, can be controlled by applying a gating potential to one of the monolayers. This opens the possibility of implementing gate-controlled $0-\pi$ transitions in Josephson junctions through the magnetic vdW bilayer weak links.
@@ I Introduction
The proximity effect between different materials modifies their properties near the interface. In the case of a superconductor/normal metal (S/N) interface, the Cooper pairs can penetrate into the normal metal at a typical distance of the order of the normal coherence length $\xi_{N}$. In the diffusive case this length is of the order of $\xi_{N}\sim\sqrt{D/2\pi T}$, where $D$ is the diffusion constant and $T$ is the temperature. In the ballistic case $\xi_{N}\sim v_{F}/2\pi T$, where $v_{F}$ is the electron Fermi velocity in the normal metal. At superconductor/ferromagnet (S/F) interfaces the singlet Cooper pairs penetrating into the ferromagnetic region are partially converted into triplet ones and acquire finite momentum. This is caused by the presence of the exchange field $h$ in the ferromagnet. Because of this field, the energy of a spin-up electron decreases by $h$, while the energy of a spin-down electron increases by $h$. To compensate this energy difference, the kinetic energy of the spin-up (down) electron increases (decreases), what gives rise to the difference of the absolute values of their momenta [1, 2, 3]. Due to the finite momentum the decay of the singlet and triplet pairs into the ferromagnet is superimposed by oscillations. The finite momentum of the pair and the resulting spatial oscillations of its wave function are the key features of the finite-momentum inhomogeneous superconductivity, predicted earlier by Fulde and Ferrell [4] and Larkin and Ovchinnikov[5]. Thus, the proximity-induced superconducting correlations in ferromagnets are often called the mesoscopic Fulde-Ferrel-Larkin-Ovchinnikov (FFLO) state [2] in the literature.
In the ballistic case the scale of the spatial oscillations of the proximity-induced pair correlations in the ferromagnet is $\sim v_{F}/2h$ if $h\gg T$ [6, 7]. In the diffusive case the penetration length and the period of oscillations are of the same order of the magnetic coherence length $\xi_{F}\sim\sqrt{D/h}\ll\xi_{N}$. The presence of inhomogeneous magnetization in the interlayer region essentially modifies the triplet proximity effect in ferromagnetic interlayers of JJs giving rise to the long-range range triplet component, which does not oscillate and decays at $\xi_{N}$ [3, 8, 9].
@@
The oscillations of the proximity-induced singlet superconducting correlations in the ferromagnet lead to $0-\pi$ transitions of the ground state superconducting phase difference in superconductor/ferromagnet/superconductor Josephson junctions (S/F/S JJs) upon varying the length of the ferromagnetic weak link [10, 2]. Spatial oscillations of proximity-induced triplet superconductivity were also predicted at superconductor/antiferromagnet (S/AF) interfaces [11]. Yet the physical mechanism is different from S/F heterostructures and does not lead to oscillations of the singlet component. Thus, it does not result in $0-\pi$ transitions in S/AF/S JJs. At the same time, $0-\pi$ transitions in S/AF/S JJs upon atomic variations of the AF interlayer length were predicted [12, 13, 14]. The physical reason for such $0-\pi$ transitions is related to the reconstruction of the Andreev bound states spectra inside the interlayer and thus atomically sharp S/AF interfaces are required to observe the effect.
@@
JJs with the $\pi$ ground state phase difference ($\pi$-JJs) are implemented as $\pi$-phase shifters in superconducting and quantum electronics [15, 16, 17]. The broad application prospects motivated an active search for novel principles and possibilities of external control of the switching between the 00 and $\pi$ ground states. Various suggestions of externally controlled $0-\pi$ transitions in Josephson junctions include the temperature induced $0-\pi$ transitions [18], $0-\pi$ transitions induced by the electrostatic gating [19, 20] in quantum dot JJs, as well as by spin-independent [21, 22, 23, 24, 25, 26, 27, 28] and spin-dependent [29, 30] nonequilibrium quasiparticle distribution, via the manipulation of the mutual orientation of the ferromagnets of a spin-valve inside a JJ [31] and by controlling the exchange field orientation in S/F/S JJs with spin-orbit coupling [32, 33]. In a two-dimensional electron gas (2DEG) $0-\pi$ transitions can also be controlled by an applied magnetic field or gating [34]. The supercurrent-induced long-range triplet component in Josephson junctions [35] provides a possibility of a supercurrent-controllable $0-\pi$ transition [36].
@@
In this paper we consider the penetration of proximity-induced superconducting correlations into bilayer magnetic vdW materials or heterostructures. It is demonstrated that at low electron transparency (interlayer hopping) between the interacting vdW layers composing the bilayer the proximity-induced Cooper pairs can be divided into local and nonlocal ones. The penetration of the nonlocal pairs into nonsuperconducting materials substantially differs from the penetration of the local pairs. In particular, the length of spatial oscillations of such pairs can be tuned by applying the gating potential to one of the vdW layers. It provides a way to identify the contribution of such pairs into the Josephson current and to generate $0-\pi$ transitions via gating. It is worth noting that the possibility to control $0-\pi$ transitions via gating results from the unique sensitivity of spatial behavior of the nonlocal pairs to this parameter and is not related to any tuning of the electron density as in experiments with quantum dots and 2DEG [19, 20, 34]. Further we demonstrate that at higher values of the interlayer hopping elements there is a region of strong hybridization between the local and nonlocal pairs, where the physical properties of all pairs differ from the conventional bulk behavior and the corresponding oscillation periods can be controlled by gating, which again gives rise to unusual and tunable behavior of the Josephson current. We investigate and compare manifestations of the nonlocal proximity-induced superconducting pairs in magnetic bilayers with antiferromagnetic, ferrimagnetic and ferromagnetic interlayer ordering. We also note that although proximity effects in vdW and layered S/F, S/F/S, F/S/F and S/F/F heterostructures with atomic thicknesses are being studied very actively [48, 49, 50, 51, 52, 44, 53, 54, 55, 56, 57, 58, 59, 60, 61], the majority of those works consider a sandwich geometry, where the system is translational invariant along the interface and contains a small number of atomic layers, which does not allow one to observe a spatial dependence of the superconducting properties. In contrast to the previous studies, we focus on the spatial behavior of the proximity-induced superconducting correlations in the plane of the layers and its dependence on the applied gating potential.
The paper is organized as follows. In Sec. II the system under consideration is described. Sec. III is devoted to analytical investigation of different proximity-induced Cooper pairs and their penetration into non-superconducting bilayers: in Sec. III.1 the generalized quasiclassical Eilenberger equation for the description of bilayer vdW heterostructures is derived and in Sec. III.2 this equation is applied to the analytical description of proximity-induced Cooper pairs. In Sec. IV results for the Josephson current through magnetic vdW bilayers are presented: in Sec. IV.1 the contribution of the nonlocal pairs in JJs with AF bilayer weak links is discussed; Sec. IV.2 demonstrates how the nonlocal pair contribution gives rises to $0-\pi$ transitions induced by gating and in Sec. IV.3 we consider how the nonlocal pairs manifest themselves in JJs with ferromagnet/normal metal interlayers. Our conclusions are formulated in Sec. V.
@@ II System and model
Figure 1: Sketch of the considered JJ. Superconducting sites are depicted by blue, while red and orange layers represent the upper ferromagnetic layer (F1) and bottom ferromagnetic layer (F2), respectively. The interlayer magnetic ordering can be of different types: AF, F or ferrimagnetic, see text for more details. Different hopping elements are also shown.
The system under consideration is shown in Fig. 1. It represents a planar JJ made of layered superconducting leads and a bilayer magnetic weak link. The system is modelled by the following tight-binding Hamiltonian on a square lattice:
@@
Here $c_{\bm{i},\sigma}$ is an annihilation operator for electrons at site $\bm{i}$ and for spin $\sigma=\uparrow,\downarrow$. $\mu_{\bm{i}}$ is the on-site energy, which is assumed to be $\mu_{\bm{i}}=\mu_{1,2}$ at sites belonging to the upper (F1) and lower (F2) ferromagnetic layers, respectively, and $\mu_{\bm{i}}=\mu_{S}$ at the sites belonging to the superconducting leads. The on-site energy determines the filling factors of the conduction band of the materials and in the case of isolated layers correspond to their chemical potentials. We assume only nearest-neighbor hopping with $t_{\bm{i}\bm{j}}=t_{x}^{F}$ being the intralayer hopping element for both ferromagnetic layers and $t_{\bm{i}\bm{j}}=t_{z}^{F}$ being the hopping element between the F1 and F2 layers. Anisotropic intralayer (interlayer) hopping elements between nearest superconducting sites are denoted by $t_{\bm{i}\bm{j}}=t_{x(z)}^{S}$ and $t_{\bm{i}\bm{j}}=t^{SF}$ is the hopping element between the upper ferromagnetic layer and the superconducting leads. $\langle\bm{i}\bm{j}\rangle$ means summation over nearest neighbors. $h.c.$ means hermitian conjugate.
$\bm{h}_{\bm{i}}$ is the exchange field in the ferromagnetic layers, which is equal to $\bm{h}_{1,2}$ in the F1 and F2 layers, respectively and is assumed to be spatially homogeneous in the plane of each layer. In superconducting leads $\bm{h}_{\bm{i}}=0$. Below we consider different physical cases: $\bm{h}_{1}=\bm{h}_{2}$ (ferromagnetic interlayer ordering), $\bm{h}_{1}=-\bm{h}_{2}$ (antiferromagnetic interlayer ordering), $\bm{h}_{1}$ is opposite to $\bm{h}_{2}$ and $|\bm{h}_{1}|\neq|\bm{h}_{2}|$ (ferrimagnetic interlayer ordering). The case $\bm{h}_{1}\neq 0$ and $\bm{h}_{2}=0$ modeling a bilayer consisting of a ferromagnetic and a paramagnetic layers is also considered. Such a bilayer structure can be implemented by exfoliation of vdW magnetic materials [62, 63, 64] or using chemical vapor deposition (CVD) processes. Thin films of some intermetallic materials, such as $\mathrm{GdIr_{2}Si_{2}}$ or $\mathrm{TbRh_{2}Si_{2}}$, seem to be good candidates [65, 66]. Also the bilayers can be assembled from magnetic and nonmagnetic monolayers of different vdW materials [42], possibly with an insulating vdW layer between them in order to exclude the direct exchange interaction and to control the magnetization orientations in the layers by the applied magnetic field.
@@ III Penetration of nonlocal correlations into nonsuperconducting regions
Here we consider a non-superconducting bilayer, which has an interface with a superconductor and investigate the proximity penetration of the Cooper pairs from the superconductor into the depth of the bilayer. Details of the geometry of the interface do not matter in this consideration. The superconductor can be located on top of a part of the bilayer (a half of the sketch presented in Fig. 1 without the second lead) or the superconductor can have a contact with both non-superconducting layers. These details only influence the amplitudes of the different types of Cooper pairs induced in the bilayer by the proximity effect, yet do not affect their spatial behavior inside the bilayer, penetration depth and oscillating period. The problem can be studied analytically in the framework of the Eilenberger equation for the quasiclassical Green’s function describing electronic behavior in the bilayer.
III.1 Derivation of the two-layer quasiclassical eilenberger equation
At first we describe the main steps of the derivation of the quasiclassical Eilenberger equation. We begin with the derivation of the Gor’kov equation for the full quantum mechanical Green’s function. Let us consider a general bilayer system [not the particular planar JJ described by the Hamiltonian (1)] described by the following tight-binding Hamiltonian:
Here $\hat{c}_{\bm{i},\sigma}=(c_{\bm{i},\sigma}^{1},c_{\bm{i},\sigma}^{2})^{T}$ is a vector composed of annihilation operators for electrons belonging to the first and the second layers at site $\bm{i}$ in the plane of each layer and for spin $\sigma=\uparrow,\downarrow$. $\hat{\mu}=\left(\begin{matrix}\mu_{1}&0\\ 0&\mu_{2}\\ \end{matrix}\right)$ is the matrix composed of on-site energies of the first and the second layers, respectively. $t$ is the hopping element in plane of each layer and $t_{12}$ is the hopping element between the first and the second layers. $\bm{\hat{h}}=\left(\begin{matrix}\bm{h}_{1}&0\\ 0&\bm{h}_{2}\\ \end{matrix}\right)$ is the matrix containing the exchange field of both layers and $\hat{\Delta}_{\bm{i}}=\left(\begin{matrix}\Delta_{1,\bm{i}}&0\\ 0&\Delta_{2,\bm{i}}\\ \end{matrix}\right)$ is the matrix composed of superconducting order parameters of the layers, which vanishes in the nonsuperconducting layer.
The Matsubara Green’s function in the two-layer formalism is an $8\times 8$ matrix in the direct product of spin, particle-hole and layer spaces. Introducing the two-layer Nambu spinor $\check{\psi}_{\bm{i}}=(\hat{c}_{{\bm{i}},\uparrow}^{1},\hat{c}_{\bm{i},\downarrow}^{1},\hat{c}_{\bm{i},\uparrow}^{2},\hat{c}_{\bm{i},\downarrow}^{2},\hat{c}_{\bm{i},\uparrow}^{1\dagger},\hat{c}_{\bm{i},\downarrow}^{1\dagger},\hat{c}_{\bm{i},\uparrow}^{2\dagger},\hat{c}_{\bm{i},\downarrow}^{2\dagger})^{T}$ we define the Green’s function as follows:
where $\langle T_{\tau}...\rangle$ means imaginary time-ordered thermal averaging. Introducing Pauli matrices in spin, particle-hole and layer spaces, as $\sigma_{k}$, $\tau_{k}$ and $\rho_{k}$ ($k=0,x,y,z$) and operator $\hat{j}$ as
$\hat{j}\hat{c}_{\bm{i}}=\sum\limits_{<\bm{i}\bm{j}>}t\hat{c}_{\bm{j}}$ (4)
and expanding $\check{G}_{\bm{i}\bm{j}}(\tau_{1}-\tau_{2})=T\sum\limits_{\omega_{m}}e^{-i\omega_{m}(\tau_{1}-\tau_{2})}\check{G}_{\bm{i}\bm{j}}(\omega_{m})$ over Matsubara frequencies $\omega_{m}=\pi T(2m+1)$, one can obtain the Gor’kov equation for Green’s function $\check{G}_{\bm{i}\bm{j}}(\omega_{m})$. The derivation is similar to that described in Ref. 61. The resulting Gor’kov equation takes the form:
$G_{\bm{i}}^{-1}\check{G}_{\bm{i}\bm{j}}(\omega_{m})=\delta_{\bm{i}\bm{j}},$ (5)
$G_{\bm{i}}^{-1}=\tau_{z}\left(\hat{j}+\hat{\mu}-\check{\Delta}_{\bm{i}}i\sigma_{y}-\bm{\hat{h}}\check{\bm{\sigma}}+t_{12}\rho_{x}\right)+i\omega_{m}.$ (6)
where $\bm{R}=(\bm{i}+\bm{j})/2$ and the integration is over $\bm{r}=\bm{i}-\bm{j}$. To further simplify calculations and to present the Gor’kov equation in a more common form we define the following transformed Green’s function:
where and below subscript $\tau$ means that the explicit matrix structure corresponds to the particle-hole space. In the quasiclassical approximation we disregard terms of the order of $1/(\xi_{S}p_{F})\ll 1$ with respect to unity, where $p_{F}$ is the Fermi momentum, $\xi_{S}=v_{F}/|\Delta|$ is the superconducting coherence length and $\hbar=1$. For this reason we can approximate $\hat{\Delta}_{\bm{i}}\approx\hat{\Delta}(\bm{R})$. The neglected terms $\sim(\bm{\nabla}\hat{\Delta}_{\bm{i}})\bm{r}\sim|\Delta_{1,2}|/(p_{F}\xi_{S})$. In general, the same is valid for the exchange field if it is inhomogeneous in the plane of the layer. With these approximations we obtain:
$\check{\tilde{G}}(\bm{R},\bm{p})=1$ (13)
Subtracting Eqs. (13) and (14) we come to the following equation:
$F([\hat{j},\check{\hat{G}}])=0$ (15)
We introduce a quasiclassical $\xi$-integrated Green’s function:
[…]
where $a$ is the lattice constant and the integration is taken over the electron energy $\xi=-2t[\cos(p_{x}a)+\cos(p_{y}a)]$ around the Fermi surface. The resulting quasiclassical Green’s function $\hat{g}$ only depends on the center of mass coordinate $\bm{R}=(\bm{i}+\bm{j})/2$ of the paired electrons with momenta $\bm{p}_{F}$ and $-\bm{p}_{F}$ at the Fermi surface and the direction of the trajectory determined by $\bm{p}_{F}$. After the $\xi$-integration of Eq. (15) the Eilenberger equation for the quasiclassical Green’s function takes the form:
$+i\bm{v}_{F}\bm{\nabla}\check{g}(\bm{R},\bm{p}_{F})=0,$ (17)
$\check{g}^{2}(\bm{R},\bm{p}_{F})=1.$ (18)
III.2 Proximity-induced superconducting correlations in a non-superconducting bilayer
Let us apply the derived Eilenberger equation to the non-superconducting bilayer, which serves as an interlayer of the considered JJ. Then the order parameter $\Delta_{\bm{i}}=0$ in both layers, $t_{12}\equiv t_{z}^{F}$ according to the Hamiltonian (1) and the generalized Eilenberger equation takes the form: